# Generalized Recurrence Criteria for Classes of Open Quantum Walks

### v1

@@ 4 A General Recurrence Criteria for Open Quantum Walks
Consider a coin $(L,R)$ of dimension $n<\infty.$ By [5], we can obtain a decomposition of $\mathfrak{h}$ in the following way:
$\mathfrak{h}=\mathfrak{R}\oplus\mathfrak{T},$
where
$\mathfrak{R}=\sup\{\mbox{supp}(\omega)\;|\;\omega\mbox{ is an invariant state for }\mathcal{L}\}.$
@@
Further, we can write
$\mathfrak{R}=\bigoplus_{\alpha\in A}\mathfrak{X}_{\alpha}=\bigoplus_{\alpha\in A}\bigoplus_{\beta\in I_{\alpha}}\mathfrak{V}_{\alpha,\beta},$
where every $\mathfrak{X}_{\alpha}$ either is a minimal enclosure or assumes a decomposition (not unique) as the sum of mutually orthogonal isomorphic minimal enclosures $\{\mathfrak{V}_{\alpha,\beta},\;\beta\in I_{\alpha}\}$. Given a minimal enclosure $\mathfrak{V}_{\alpha},$ let $\tau_{\alpha}$ be the unique invariant state with support on $\mathfrak{V}_{\alpha}.$ Introducing the parameter
$m_{\alpha}=\mathrm{Tr}\left(L\tau_{\alpha}L^{*}\right)-\mathrm{Tr}\left(R\tau_{\alpha}R^{*}\right),$
we have by [5, Lema 4.2] that $m_{\alpha}$ is independent of the particular minimal enclosure $\mathfrak{V}_{\alpha}$ in $\mathfrak{X}_{\alpha}.$ Therefore, we can combine Proposition 3.4, Theorem 3.7 and the discussion above to state the property
@@
  • $\textmd{Tr}\left(L\tau_{\alpha}L^{*}\right)=\dfrac{1}{2}\Rightarrow$ the walk is $\tau_{\alpha}$-recurrent for all $\tau\in\mathcal{S}(\mathcal{Y}_{\alpha})$;
  • $\textmd{Tr}\left(L\tau_{\alpha}L^{*}\right)\neq\dfrac{1}{2}\Rightarrow$ the walk is $\tau_{\alpha}$-transient for all $\tau\in\mathcal{S}(\mathcal{Y}_{\alpha}).$
@@
The following result is due to [5] and will be crucial to our criteria. To state it, consider a coin $(L,R),$ and let $\mathcal{Y}$ be an enclosure for $\mathcal{L}.$ Denoting $p_{\mathcal{Y}}$ the orthogonal projection onto $\mathcal{Y},$ we can define the associated absorption operator (see [4]):
@@
Consider a coin $(L,R)$ with the decomposition of $\mathfrak{h}$ given on Equation (3), and let $a_{\alpha}(\rho)=\mathbb{E}_{\rho}\left[\mathrm{Tr}\left(A(\mathcal{Y}_{\alpha})\rho_{0}\right)\right].$ Then
@@
Let us start an OQW induced by a finite-dimensional coin $(L,R)$ and we assume that, in the $n$-th step, the density operator is $\rho_{n}.$ Note that each $\mathcal{Y}_{\alpha}$ on the Proposition is an enclosure, and any $\rho_{n}\in\mathcal{S}(\mathfrak{h})$ can be written in the form
$\rho_{n}=\sum_{\alpha\in A}a_{\alpha}^{(n)}\tau_{\alpha}^{(n)}+\tau^{(n)},$
where $\tau_{\alpha}^{(n)}$ is a positive operator with support on $\mathcal{Y}_{\alpha}$ and $\tau^{(n)}$ is either null or a positive operator with support on $\mathcal{X}$ as in Equation (3). There are three possibilities concerning recurrence when $\tau^{(0)}$ is non-null and its support is not disjoint from $\mathcal{X}.$
1. $(L,R)$ is recurrent for $\mathcal{Y}_{\alpha}$ for some $\alpha$ and supp$\left(\mathcal{L}(\rho)\right)\cap\mathcal{Y}_{\alpha}\neq\emptyset$: in this case, the walk is $\rho$-recurrent. Indeed, concerning the quantum trajectories $(\rho_{n})_{n}$, we could refresh the walk on some instant $n$ at some vertex $\lvert X_{n}\rangle,$ which is $\rho_{n}$-recurrent, since supp$\left(\rho_{n}\right)\cap\mathcal{Y}_{\alpha}\neq\emptyset$ for a sufficient large $n$. If there was some density $\sigma$ with supp$(\sigma)\cap\mathcal{X}$ and $\sigma$-transient, there would be some enclosure contained on $\mathcal{X},$ which is a contradiction by Proposition 3.4. The conclusion is held by the homogeneity of the coin.
2. $(L,R)$ is transient for all $\mathcal{Y}_{\alpha}$ satisfying supp$\left(\mathcal{L}(\rho)\right)\cap\mathcal{Y}_{\alpha}\neq\emptyset$: the walk is $\sigma$-transient for all $\sigma\in\mathcal{S}(\mathcal{X})$. Indeed, in this case, starting the walk on vertex $X_{0}=0,$ $X_{n}/n$ converges (when $n\rightarrow\infty$) to
@@
as given by Equation (4), where the second equality follows from $m_{\alpha}\neq 0,$ since $(L,R)$ is $\mathcal{Y}_{\alpha}$ transient. Thus, $X_{n}$ converges either to $\infty$- ∞ or $\infty$,∞ , in probability, imposing that the walk is $\rho$-transient and thus $(L,R)$ is transient on $\mathcal{S}(\mathcal{X}).$
@@
Consider a coin $(L,R)$ of dimension $d<\infty.$ Consider the decomposition of $\mathfrak{h}$ as given in Equation (3), and $\tau_{1},\ldots,\tau_{m}$ the respectively invariant states of $\mathcal{L}$ in $\mathcal{Y}_{\alpha},\alpha=1,\ldots,m.$
@@
$(2)\quad\mathrm{Tr}(L\tau_{j_{k}}L^{*})=\dfrac{1}{2}\;$ for a set of indices $j_{k}\in J\subset\{1,\ldots,m\}$ and $\mathrm{Tr}(L\tau_{j_{l}}L^{*})\neq\dfrac{1}{2}\;$ for the remaining indices $j_{l}\in J^{C},$ and $\sigma$ is a density such that supp$(\sigma)\cap\mathcal{X}\neq\emptyset.$
(2.1) supp$\left(\mathcal{L}(\sigma)\right)\cap\mathcal{Y}_{\alpha}\neq\emptyset$ for some $\alpha\in J,$ then (L,R) is
$\begin{cases}\rho\text{-transient for }\rho\in\mathcal{D}_{1}:=\mathcal{S}(\oplus_{k\in J^{C}}\mathcal{Y}_{k})\\ \rho\text{-recurrent for }\rho\in\mathcal{D}_{1}^{C}.\end{cases}$
(2.2) supp$\left(\mathcal{L}(\rho)\right)\cap\mathcal{Y}_{\alpha}=\emptyset$ for all $\alpha\in J,$ then (L,R) is
$\begin{cases}\rho\text{-transient for }\rho\in\mathcal{D}_{2}:=\mathcal{S}(\oplus_{k\in J^{C}}\mathcal{Y}_{k}\oplus\mathcal{X})\\ \rho\text{-recurrent for }\mathcal{D}_{2}^{C}\end{cases}$
@@ 5 Examples
Our first examples will consider lazy OQWs of dimension 2.
@@
Consider the Lazy OQW induced by the following coin of dimension 3: $(L,B,R),$ where
@@
Let us consider the coin given in [5, Section 6.2]:
@@
Consider an OQW induced by the coin $(L,R),$ where
@@ 6 Conclusion
We present new general recurrence criteria for classes of homogeneous OQWs, whose dynamics are governed by Kraus operators that are induced by a finite-dimensional coin.
In the context of these OQWs, we presented three distinct criteria for analyzing recurrence. Firstly, Theorem 3.7 introduces a criterion for ergodic Lazy coins, where the recurrence is determined by the operators $L$ and $R$ and by the only invariant state $\rho_{\infty}$ associated with the corresponding auxiliary map.
Next, Theorem 3.9 focuses on Lazy OQWs of dimension 2, offering a complete and easy-to-apply criterion. The simplicity of this criterion comes from the fact that it depends on the knowledge of the eigenvectors of the operators $L$, $B$, and $R$, allowing a practical analysis of the recurrence of these processes that are not necessarily ergodic.
Finally, Theorem 4.3 extends the analysis to non-lazy coins, using a decomposition of the vector space where the operators $L$, $B$, and $R$ act. By checking recurrence in vector subspaces that contain unique invariant states, this general criterion is extended to any initial density, significantly expanding the scope of recurrence analysis for OQWs in more complex configurations, since it is valid for any finite dimensional coin.

### v2

@@ 4 A General Recurrence Criteria for Open Quantum Walks
Following the approach described in [6, Section 4], as well as the references therein, we can obtain for a coin $(L,R)$ a decomposition
$\mathfrak{h}=\mathfrak{R}\oplus\mathfrak{T},\qquad\mathfrak{R}=\sup\{\mbox{supp}(\omega)\;|\;\omega\mbox{ is an invariant state for }\mathcal{L}\}.$
@@
Further, every $\mathfrak{X}_{\alpha}$ either is a minimal enclosure or assumes a decomposition (not unique) as the sum of mutually orthogonal isomorphic minimal enclosures $\{\mathfrak{V}_{\alpha,\beta},\;\beta\in I_{\alpha}\}:$
$\mathfrak{X}_{\alpha}=\bigoplus_{\beta\in I_{\alpha}}\mathfrak{V}_{\alpha,\beta}.$
Given a minimal enclosure $\mathfrak{V}_{\beta},$ let $\tau_{\beta}$ be the unique invariant state with support on $\mathfrak{V}_{\beta}.$ Introducing the parameter
$m_{\beta}=\mathrm{Tr}\left(L\tau_{\beta}L^{*}\right)-\mathrm{Tr}\left(R\tau_{\beta}R^{*}\right),$
we have by [6, Lemma 4.2] that $m_{\beta}$ is independent of the particular minimal enclosure $\mathfrak{V}_{\beta}$ in $\mathfrak{X}_{\alpha}.$ Therefore, we can combine Proposition 3.4, Theorem 3.7 and the discussion above to state the property
@@
  • $\textmd{Tr}\left(L\tau_{\alpha}L^{*}\right)=\dfrac{1}{2}\Rightarrow$ the walk is $\tau_{\alpha}$-recurrent for all $\tau\in\mathcal{D}(\mathcal{Y}_{\alpha})$;
  • $\textmd{Tr}\left(L\tau_{\alpha}L^{*}\right)\neq\dfrac{1}{2}\Rightarrow$ the walk is $\tau_{\alpha}$-transient for all $\tau\in\mathcal{D}(\mathcal{Y}_{\alpha}).$
@@
The following result is due to [6] and will be crucial to our last criterion. To state it, consider a coin $(L,R),$ and let $\mathcal{Y}$ be an enclosure for $\mathcal{L}.$ Denoting $p_{\mathcal{Y}}$ the orthogonal projection onto $\mathcal{Y},$ we can define the associated absorption operator (see [5]):
@@
Consider a coin $(L,R)$ with the decomposition of $\mathfrak{h}$ given on Equation (5), and let $a_{\alpha}(\rho)=\mathbb{E}_{\rho}\left[\mathrm{Tr}\left(A(\mathcal{Y}_{\alpha})\rho_{0}\right)\right].$ Then we have the convergence in law
@@
We remark that the convergence in law brought by corollary 4.2 can be upgraded to a convergence a.s., as demonstrated in [9, Theorem 3.5.2].
Let us start a HOQW induced by a coin $(L,R).$ There are two possibilities concerning recurrence when $\text{supp}(\rho_{0})\not\perp\mathcal{X}:$
1. $(L,R)$ is recurrent for $\mathcal{Y}_{\alpha}$ for some $\alpha$ and supp$\left(\mathcal{L}^{n}(\rho_{0})\right)\not\perp\mathcal{Y}_{\alpha}$ for some $n\in\mathbb{N}$: in this case, the walk is $\rho_{0}$-recurrent. Indeed, concerning the quantum trajectories $(\rho_{n})_{n}$, we could refresh the walk on some instant $n$ at some vertex $\lvert X_{n}\rangle,$ which is $\rho_{n}$-recurrent, since supp$\left(\mathcal{L}^{n}(\rho_{0})\right)\not\perp\mathcal{Y}_{\alpha}$ for a sufficiently large $n$. If there was some density $\sigma$ with supp$(\sigma)\not\perp\mathcal{X}$ and $\sigma$-transient, there would be some enclosure not orthogonal to $\mathcal{X},$ which is a contradiction by Proposition 3.4. The conclusion is held by the homogeneity of the coin, since, in this case, the walk either is $\rho_{s}$-recurrent for every $s=0,1,2,\ldots$ or it is $\rho_{s}$-transient for every $s=0,1,2,\ldots$, analogously to the proof of Proposition 3.4.
2. $(L,R)$ is transient for all $\mathcal{Y}_{\alpha}$ satisfying supp$\left(\mathcal{L}^{n}(\rho_{0})\right)\not\perp\mathcal{Y}_{\alpha}$ for every $n\in\mathbb{N}$: the walk is $\sigma$-transient for all $\sigma\in\mathcal{D}(\mathcal{X})$. Indeed, in this case, starting the walk on vertex $X_{0}=0,$ $X_{n}/n$ converges (when $n\rightarrow\infty$) to
@@
as given by Equation (6), where the second equality follows from $m_{\alpha}\neq 0,$ since $(L,R)$ is $\mathcal{Y}_{\alpha}$ transient. Thus, $X_{n}$ converges either to $\infty$- ∞ or $\infty$,∞ , in probability, imposing that the walk is $\rho$-transient and thus $(L,R)$ is transient on $\mathcal{D}(\mathcal{X}).$
@@
Consider a coin $(L,R)$ of dimension $d.$ Consider the decomposition of $\mathfrak{h}$ as given in Equation (5), and $\tau_{1},\ldots,\tau_{m}$ the respectively invariant states of $\mathcal{L}$ in $\mathcal{Y}_{\alpha},\alpha=1,\ldots,m.$
@@
$(2)\quad\mathrm{Tr}(L\tau_{j_{k}}L^{*})=\dfrac{1}{2}\;$ for a set of indices $j_{k}\in J\subset\{1,\ldots,m\}$ and $\mathrm{Tr}(L\tau_{j_{l}}L^{*})\neq\dfrac{1}{2}\;$ for the remaining indices $j_{l}\in J^{C},$ and $\sigma$ is a density such that supp$(\sigma)\not\perp\mathcal{X}.$
(2.1) supp$\left(\mathcal{L}^{n}(\sigma)\right)\not\perp\mathcal{Y}_{\alpha}$ for some $\alpha\in J$ and some $n\in\mathbb{N},$ then $(L,R)$ is
$\begin{cases}\rho\text{-transient for }\rho\in\mathcal{D}_{1}:=\mathcal{D}(\oplus_{k\in J^{C}}\mathcal{Y}_{k})\\ \rho\text{-recurrent for }\rho\in\mathcal{D}_{1}^{C}.\end{cases}$
(2.2) supp$\left(\mathcal{L}^{n}(\rho)\right)\perp\mathcal{Y}_{\alpha}$ for all $\alpha\in J$ and all $n\in\mathbb{N},$ then $(L,R)$ is
$\begin{cases}\rho\text{-transient for }\rho\in\mathcal{D}_{2}:=\mathcal{D}(\oplus_{k\in J^{C}}\mathcal{Y}_{k}\oplus\mathcal{X})\\ \rho\text{-recurrent for }\mathcal{D}_{2}^{C}\end{cases}$
@@
Remark 4.4.
It is worth observing that $\mathcal{D}_{2}$ in Theorem 4.3 actually arises from an enclosure due to the orthogonality between supp$\left(\mathcal{L}^{n}(\rho)\right)$ and $\mathcal{Y}_{\alpha}$; otherwise, it would contradict Proposition 3.4.
@@ 5 Examples
Our first examples will consider lazy HOQWs of dimension 2.
@@
Consider the Lazy HOQW induced by the following coin of dimension 3: $(L,B,R),$ where
@@
Let us consider the coin given in [6, Section 6.2]:
@@
Consider a HOQW induced by the coin $(L,R),$ where
@@ 6 Conclusion
We present new general recurrence criteria for classes of HOQWs, whose dynamics are governed by Kraus operators that are induced by a finite-dimensional coin.
In the context of these HOQWs, we presented three distinct criteria for analyzing recurrence. Firstly, Theorem 3.7 introduces a criterion for ergodic Lazy coins, where the recurrence is determined by the operators $L$ and $R$ and by the only invariant state $\rho_{\infty}$ associated with the corresponding auxiliary map.
Next, Theorem 3.9 focuses on Lazy HOQWs of dimension 2, offering a complete and easy-to-apply criterion. The simplicity of this criterion comes from the fact that it depends on the knowledge of the eigenvectors of the operators $L$, $B$, and $R$, allowing a practical analysis of the recurrence of these processes that are not necessarily ergodic.
Finally, Theorem 4.3 extends the analysis to non-lazy coins, using a decomposition of the vector space where the operators $L$, $B$, and $R$ act. By checking recurrence in vector subspaces that contain unique invariant states, this general criterion is extended to any initial density, significantly expanding the scope of recurrence analysis for HOQWs in more complex configurations, since it is valid for any finite-dimensional coin.